# Wave-driven phase wave patterns in a ring of FitzHugh-Nagumo oscillators

### v1

@@ I Introduction
$\begin{array}[]{ @{} l @{} }u_{t}=-u^{3}+u-v,\\[3.0pt] v_{t}=\varepsilon(u-bv+a).\end{array}$ (1)
@@
This study draws inspiration from biological systems, where many critical networks are described by activator-inhibitor pairs, which can be modeled by the FHN model. Our primary motivation stems from the intricate interactions between the Cyclin B-Cdk1 enzyme and the APC/C protein complex that regulate mitosis in the cell cytoplasm[8, 9, 10, 11, 12], alongside the interplay between Rho GTPase and F-Actin that governs cytoskeletal behaviors at the cell cortex[13, 14, 15]. These interactions are crucial for coordinating cell division during cytokinesis.
In this work, we utilize a dual FHN system setup to simulate the intricate dynamics of a cell’s cytoplasm and cortex. The inner system, representing the cytoplasm, exerts a unidirectional inhibitory influence on the outer system, analogous to the cortex, through pacemaker-driven waves. This setup is designed to reflect the key dynamics of the Cyclin B-Cdk1 (APC/C) and Rho GTP (F-Actin) networks. All simulations are conducted in dimensionless spatial and temporal units to maintain the model’s applicability across other contexts.
@@
Figure 1: Model consisting of two coupled FHN Systems A. The depicted system consists of a disc (internal system) encircled by a ring (external system), each governed by the FHN equation within a circular area of radius $R=50$. The internal system is divided into two distinct zones: an oscillatory Region 1 centered within a radius $r=10$ at $\bm{x}=(-25,0)$; and an adjacent excitable Region 2 (see model parameters at Tab. 1). The external system mirrors the parameters of Region 1 but incorporates a coupling factor $c$ that modulates its parameter $a(u_{i})$ based on the internal field $u_{i}$. B. The parameter space ($a,b$) delineating areas with distinct dynamical behaviors. Panels C-E further elaborate on these dynamics by presenting time series and phase space trajectories (including nullclines) for oscillatory, excitable, and non-excitable states, respectively.
@@ II Model system
The internal system functions exclusively as a driver for the external system’s dynamics, remaining unaffected by the latter’s behavior. Consequently, while the external system can show intricate dynamics stemming from its interaction with the internal system, the internal system itself exhibits quite straightforward behavior that can be precisely defined. In this setup, the oscillatory region (region 1), acting as a pacemaker, maintains a consistent oscillation period of $T_{i}=2200$. This rhythmic activity induces excitation in the adjacent region (Region 2), leading to the generation of traveling pulses. Despite the different curvatures, these pulses travel at a constant speed of $v\approx 0.059$ (Supp.Fig. S1) and have constant width of $\approx 18$[16]. This enables the precise prediction of the pulse’s arrival time at the external system’s boundary. We thus simplify the simulations by capturing the waveform of the traveling pulse and reintroducing it periodically, aligning with the anticipated timing (Fig. 2).
@@ X Data availability statement
The numerical codes to reproduce the figures in this study are openly available in GitLab[34], and as an archived repository on RDR by KU Leuven [Upcoming].

### v2

@@ I Introduction
$\begin{array}[]{ @{} l @{} }u_{t}=-u^{3}+u-v,\\[3.0pt] v_{t}=\varepsilon(u-bv+a).\\ \end{array}$ (1)
@@
This study draws inspiration from biological systems, where many regulatory networks are described by activator-inhibitor pairs, which can be modeled by the FHN model. Our primary motivation stems from the complex cell cycle regulatory system, where multiple parts exhibit activator-inhibitor dynamics We specifically focus on how two key regulatory modules collectively coordinate cell division during cytokinesis.
On the one hand, interactions between two protein complexes regulate mitosis in the cell cytoplasm: cyclin B - Cdk1, consisting of the protein cyclin B and an enzyme called cyclin-dependent-kinase 1 (Cdk1), and the Anaphase Promoting Complex / Cyclosome (APC/C) [8, 9, 10, 11, 12]. Additionally, the interplay between Rho GTPase and F-Actin governs cytoskeletal behaviors at the cell cortex [13, 14, 15]. Actin proteins bind to each other to form filaments, and such filaments, when polymerizing against a cell’s membrane, produce a force that can deform the membrane.
Recent experimental studies in starfish oocytes and embryos have shown that cortical dynamics respond to internal cytoplasmic signals, i.e., cyclin B-Cdk1 enzyme activity[16, 17]. By inhibiting one of its positive regulators, Ect2 [18], a cytosolic gradient of cyclin B-Cdk1 controls actin-driven waves of contraction [16]. However, studies in other model organisms (i.e., frog egg extracts and fly embryos) have found pacemaker-driven waves of Cdk1 activity traveling through the cytoplasm [19, 20, 21, 22].
In this work, we characterize the range of dynamical behaviors that can occur at the cell cortex when driven by a cytoplasmic wave traveling at constant speed. We consider this driving to be unidirectional, as no biochemical coupling from the cortex to the cytoplasm has been identified. While membrane deformations could, in principle, influence the cytoplasm as well, recent work demonstrated that under small deformations, these influences are negligible [23].
We utilize a dual FHN system setup to simulate the intricate dynamics of a cell’s cytoplasm and cortex. The inner system, representing the cytoplasm, exerts a unidirectional inhibitory influence on the outer system, analogous to the cortex, through pacemaker-driven waves. This setup is designed to reflect the key dynamics of the Cyclin B-Cdk1 (APC/C) and Rho GTP (F-Actin) networks. By using a system of two coupled FHN models, we focus on the generic properties of wave-driven patterning rather than aiming to describe the detailed biology. As such, this analysis could be relevant for other systems where an activator-inhibitor region is driven by a traveling wave. All simulations are conducted in dimensionless spatial and temporal units to maintain the model’s applicability across other contexts.
@@
Figure 1: Model consisting of two coupled FHN Systems A. The depicted system consists of a disc (internal system) encircled by a ring (external system), each governed by the FHN equation within a circular area of radius $R=50$. The internal system is divided into two distinct zones: an oscillatory Region 1 centered within a radius $r=10$ at $\bm{x}=(-25,0)$; and an adjacent excitable Region 2 (see model parameters at Tab. 1). The external system mirrors the parameters of Region 1 but incorporates a coupling factor $c$ that modulates its parameter $a(u_{i})$ based on the internal field $u_{i}$, as illustrated by the white dashed line and the black arrow in panel B. B. Representation in the parameter space ($a,b$) that illustrates all the areas with distinct dynamical behaviors that appear in our model. Panels C-E further elaborate on these dynamics by presenting time series and phase space trajectories (including nullclines) for oscillatory, excitable, and non-excitable states, respectively.
@@ II Model system
All the model parameters are given in Table 1, and their respective biological motivations are as follows. Both systems exhibit asymmetric spike-like oscillations ($a>0$), with the cytoplasmic system being approximately ten times slower ($\tau$). The overall system size (radius R) is set to 50 dimensionless units. For reference, a Xenopus laevis frog egg is approximately 1 millimeter in diameter. The pacemaker region (Region 1 - radius r) is chosen to be 10 dimensionless spatial units. Previous studies have demonstrated that the nucleus acts as a pacemaker to initiate waves in the early Xenopus laevis cytoplasm [20, 21, 24]. Both nuclear size and cell size change during the early cleavage stages of the frog embryo, with the nuclear-cytoplasmic ratio ranging from about 0.1% to 10% of the total cell size [25]. The period of our pacemaker region is set to approximately 2200 dimensionless time units, which corresponds to a typical cell cycle period of 30 minutes in the early Xenopus laevis embryo. By normalizing time and space in this manner, we derive realistic values for the diffusion coefficients, based on Ref. [26]. These coefficients were further fine-tuned to match the observed cytoplasmic and cortical wave speeds [19, 21, 27]. Finally, the coupling parameters were chosen to replicate the inhibitory dynamics exerted by the cytoplasm on the cortex.
The internal system functions exclusively as a driver for the external system’s dynamics, remaining unaffected by the latter’s behavior. Consequently, while the external system can show intricate dynamics stemming from its interaction with the internal system, the internal system itself exhibits quite straightforward behavior that can be precisely defined. Note that changes in the pacemaker position lead to changes in the external dynamics, but changes in the radius do not (Supp.Fig. S1).
In this setup, the oscillatory region (region 1), acting as a pacemaker, maintains a consistent oscillation period, resembling the nucleus of $T_{i}=2200$. This rhythmic activity induces excitation in the adjacent region (Region 2), leading to the generation of traveling pulses. Despite the different curvatures, these pulses travel at a constant speed of $v\approx 0.059$ (Supp.Fig. S2) and have constant width of $\approx 18$[28]. This enables the precise prediction of the pulse’s arrival time at the external system’s boundary. We thus simplify the simulations by capturing the waveform of the traveling pulse and reintroducing it periodically, aligning with the anticipated timing (Fig. 2).
@@ X Data availability statement
The numerical codes to reproduce the figures in this study are openly available in GitLab[43], and as an archived repository on RDR by KU Leuven[44].
@@ C.2 Supplementary Figures
Figure S1: Effect of the pacemaker parameters. A. Changing the position of the pacemaker affects the relative arrival timing to the driven system, thereby modifying the window of emergent dynamics. In the case of a centered pacemaker, no dynamics emerge. B. Changing the radius affects the overall timing of arrival to the driven system but not the relative timing, thus trivially leading to conserved emergent dynamics.